# Synthesizing Sheet Music Problems for Evaluation and Reinforcement Learning

### v1

@@ Abstract
Enhancing the ability of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) to interpret sheet music is a crucial step toward building AI musicians. However, current research lacks both evaluation benchmarks and training data for sheet music reasoning. To address this, we propose the idea of synthesizing sheet music problems grounded in music theory, which can serve both as evaluation benchmarks and as training data for reinforcement learning with verifiable rewards (RLVR). We introduce a data synthesis framework that generates verifiable sheet music questions in both textual and visual modalities, leading to the Synthetic Sheet Music Reasoning Benchmark (SSMR-Bench) and a complementary training set. Evaluation results on SSMR-Bench show the importance of models’ reasoning abilities in interpreting sheet music. At the same time, the poor performance of Gemini 2.5-Pro highlights the challenges that MLLMs still face in interpreting sheet music in a visual format. By leveraging synthetic data for RLVR, Qwen3-8B-Base and Qwen2.5-VL-Instruct achieve improvements on the SSMR-Bench. Besides, the trained Qwen3-8B-Base surpasses GPT-4 in overall performance on MusicTheoryBench and achieves reasoning performance comparable to GPT-4 with the strategies of Role play and Chain-of-Thought. Notably, its performance on math problems also improves relative to the original Qwen3-8B-Base. Furthermore, our results show that the enhanced reasoning ability can also facilitate music composition. In conclusion, we are the first to propose the idea of synthesizing sheet music problems based on music theory rules, and demonstrate its effectiveness not only in advancing model reasoning for sheet music understanding but also in unlocking new possibilities for AI-assisted music creation. Our code and resources are available at GitHub.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) have inspired researchers to explore the potential of developing AI musicians (Qu et al., 2025; Bradshaw & Colton, 2025; Wang et al., 2024). Given that sheet music is the universal language of musicians, the ability to read and interpret it is an essential step for AI musicians  (Yuan et al., 2024; Wang et al., 2025). This requires a grasp of the core concepts of music theory, which provide the structural and expressive framework necessary for musical interpretation. Enhancing LLMs and MLLMs to interpret sheet music involves not just recognizing musical symbols but also modeling the cognitive processes involved in understanding music.
However, research on sheet music reasoning remains limited, with a lack of standardized evaluation benchmarks and training datasets. For example, ChatMusician (Yuan et al., 2024) introduces MusicTheoryBench, a benchmark consisting of 367 human-designed questions, some of which specifically target sheet music understanding in ABC notation, a textual format of sheet music. MMMU (Yue et al., 2024) provides a small set of sheet music understanding questions in image format, though only 30 of them are publicly released with answers. In contrast, MusiXQA (Chen et al., 2025a) contains a larger set of image-based sheet music data, but it mainly targets the recognition of individual elements rather than reasoning about their connections within music theory.
@@
Inspired by recent human-programmatic approaches to data synthesis (Liu et al., 2025; Chen et al., 2025b; Li et al., 2025), we propose to leverage the rules of music theory to programmatically generate an endless supply of verifiable reasoning problems about sheet music. Such a resource is invaluable for creating robust evaluation benchmarks and for training models with verifiable rewards (Shao et al., 2024; Yu et al., 2025) to improve their reasoning abilities. The practical implementation of this concept is our novel data synthesis framework. This system is designed to generate questions with staff notation presented in both textual and visual formats and includes nine Question Template Classes, categorized into Rhythm, Chord, Interval, and Scale. We use this framework to build the Synthetic Sheet Music Reasoning Benchmark (SSMR-Bench). The benchmark includes an evaluation set with 1,600 textual and 1,600 visual QA pairs, and a larger training set with 8,000 pairs for each modality.
@@
We evaluate SSMR-Bench using a range of state-of-the-art LLMs and MLLMs. For the textual QA task, which represents sheet music in ABC notation, Qwen3-32B demonstrates performance on par with Deepseek-V3, while Deepseek-R1 exceeds Deepseek-V3 by 18.69% overall. This highlights the model’s enhanced reasoning capability in understanding sheet music. For the visual QA task, Gemini2.5-Pro achieves the best performance with an average score of 56.62%, surpassing Qwen2.5-VL-72B-Instruct, InternVL3-78B, and other models. However, this performance is significantly lower than when the same questions are presented in textual format, highlighting the challenges of interpreting sheet music from images.
We train Qwen3-8B-Base and Qwen2.5-VL-7B-Instruct on our synthetic dataset using Group Relative Policy Optimization (GRPO,  Shao et al., 2024). Both models achieve substantial improvements. Specifically, the trained Qwen3-8B-Base achieved a 13.06% improvement on our SSMR-Bench over Qwen3-8B, reaching performance comparable to GPT-4.1-mini. Similarly, the trained Qwen2.5-VL-7B-Instruct improves its ability to recognize and analyze sheet music from images, allowing it to outperform Gemini2.5-Pro. To verify the effectiveness of our synthetic data, we further evaluate the trained Qwen3-8B-Base on MusicTheoryBench (Yuan et al., 2024). The results show that the model outperforms GPT-4 in overall performance on MusicTheoryBench and achieves reasoning performance comparable to GPT-4 with Role Play and Chain-of-Thought. Besides, training Qwen3-8B-Base on our synthetic data also boosts its performance on math problems, surpassing the original model’s capabilities. Furthermore, the improved reasoning ability in sheet music understanding can also facilitate music composition.
@@
  • We are the first to propose the idea of leveraging music theory rules to programmatically synthesize verifiable sheet music problems, using them as both an evaluation benchmark and a training set for Reinforcement Learning with Verifiable Rewards (RLVR).
@@
2 Related Work
2.1 Sheet Music Understanding
The application of artificial intelligence to sheet music analysis has consistently been a focal point for the research community (Fornés et al., 2012; Shatri & Fazekas, 2024). Optical Music Recognition (OMR) employs AI-driven tools to interpret music notation, necessitating both specialized expertise and domain knowledge (Ma et al., 2024; Tuggener et al., 2018; Calvo-Zaragoza & Rizo, 2018). Some research in this area focuses on the recognition of handwritten music scores, introducing widely used benchmarks and datasets (Torras et al., 2024; Mayer et al., 2024), while other studies propose novel methodologies to address the underlying challenges (Rebelo et al., 2012; Rıos-Vila et al., 2024; R´ıos-Vila et al., 2024; Li et al., 2023).
While significant progress has been made, some researchers are expanding the study to Sheet Music Question Answering (QA), moving beyond traditional music recognition and image modalities. MMMU (Yue et al., 2024) provides music QA, highlighting the challenges of sheet music QA based on images. Meanwhile, there is growing interest in using LLMs for symbolic music (Wu et al., 2024; Wang et al., 2025; Qu et al., 2025), where sheet music is represented in textual formats, such as the widely used ABC notation.  Yuan et al. (2024) introduces a MusicTheoryBench that includes knowledge and reasoning categories to evaluate the music understanding abilities of LLMs. However, the manually crafted nature of the two benchmarks hinders data scalability, which is the problem we aim to address in this work.
2.2 Reinforcement Learning Through Synthetic Data
@@ 3 Synthesize Music Reasoning Questions
3.1 Data Synthesize Framework
Figure 1: Overview of synthesizing verifiable sheet music questions. Experts define rule-based Question Template Classes that generate questions and answers from ABC notation, where the music sheet can be converted into an image, enabling the construction of VQA-style questions. The process is fully programmatic, without LLMs.
Music is traditionally written on a staff, a set of five lines where visual symbols represent the core elements of music theory. The vertical position of a notehead indicates its pitch, while its shape (e.g., hollow or filled, with a stem or flag) defines its rhythm, or duration. These visual cues are organized into measures, separated by barlines, according to a time signature that dictates the underlying rhythmic pulse. While staff notation is intuitive for human musicians, its graphical nature is cumbersome for programmatic analysis. To bridge this gap, we utilize ABC notation, a text-based format that encodes the same musical information in a machine-readable way. In ABC, pitches are represented by letters (C, D, E), rhythmic values are specified with numerical modifiers, and measures are delineated by vertical bars (|). This format allows us to translate the abstract rules of music theory into parsable strings. Examples of ABC notation are provided in Appendix A.
Our framework is designed to generate questions that test a model’s understanding of these fundamental principles, from rhythmic calculation and chord identification to the analysis of intervals and scales. We outline the process of synthesizing sheet music questions based on music theory rules in Figure 1. Sheet music is first obtained in ABC notation. Then, the experts design specialized Question Template Classes derived from music theory rules, each capable of generating both questions and corresponding answers from the ABC representation. Applying these templates produces textual QA pairs. Subsequently, the sheet music in each question can be converted into images using abcm2ps111abcm2ps converts ABC notation to PostScript sheet music. https://github.com/lewdlime/abcm2ps and ImageMagick [^2] Tool for converting images, e.g., PS or SVG to JPG. https://github.com/ImageMagick/ImageMagick, enabling the construction of VQA questions. The entire pipeline is fully programmatic, without reliance on LLMs, ensuring controllability and interpretability.
@@ 3.2 Data Construction
The scalable synthesis framework we designed currently provides 9 types of questions. All questions are categorized into four categories: Rhythm, Chord, Interval, and Scale, which are the fundamental elements of sheet music. The cases for each question type are provided in Appendix A.2.
  • Rhythm Questions challenge models to understand the rhythm of sheet music. These questions require models to calculate note values and analyze the music from a holistic perspective.
To solve these questions, LLMs and MLLMs need basic music knowledge, sheet music reading skills, and reasoning ability.
Through our data synthesis framework, we introduce the Synthesize Sheet Music Reasoning Benchmark (SSMR-Bench), comprising 1,600 textual and 1,600 visual questions. The sheet music used to generate questions is sampled from MelodyHub (Wu et al., 2024), a large-scale dataset in ABC notation. The distribution of question types and categories within the benchmark is illustrated in Figure 2. Furthermore, we offer a comprehensive training set, consisting of 8,000 textual and 8,000 visual questions. A comparative analysis of our framework with prior music sheet-related benchmarks is presented in Table 1. Our framework produces both visual and textual modality questions for music sheet reasoning and constructs a verifiable training set without human curation, thereby enabling scalable data generation.
Table 1: Comparison of music-related QA datasets across different aspects. Checkmarks (✓) denote presence, crosses (✗) absence, and triangles ($\triangle$) partial coverage.
@@
MusicTheoryBench (Yuan et al., 2024) $\triangle$
Ours
@@ 4.1 Experimental Setup
RL Practice.
We employ Qwen3-8B-Base for reinforcement learning in the textual modality and Qwen2.5-VL-7B-Instruct for the visual modality. In addition, we include a variant of Qwen3-8B-Base trained on 8,000 math problems sampled from OpenR1 (Hugging Face, 2025) to compare against the effectiveness of synthetic music reasoning data. Training is performed using the GRPO algorithm with a batch size of 128, a rollout number of 8, and an update batch size of 64. The KL loss and entropy loss are disabled by setting their coefficients to zero. The reward function relies solely on accuracy, without incorporating format-related signals. The training template is provided in Appendix B. The total number of training steps is 300.
@@ 4.2 Results on Synthetic Music Reasoning Benchmark
Results on Textual Modality.
The results for textual QA show that the LLMs, such as Deepseek-R1, outperform others significantly across all categories. Specifically, Deepseek-R1 achieves the highest overall accuracy of 93.63%, with exceptional performance in Scale (99.75%). In contrast, models like Qwen2.5-32B-Instruct and Llama3.1-70B-Instruct exhibit lower accuracies, with scores of 51.25% and 43.94%, respectively. Llama3.1-8B-Instruct shows the lowest performance with an overall score of 20.94%.
Results on Visual Modality.
Comparison Between Modality.
The lower performance of MLLMs in the visual modality can be attributed to the unique challenges posed by music element recognition. Unlike textual data, which is relatively machine-friendly, musical staff images contain intricate visual patterns. To interpret these images, MLLMs must first recognize diverse musical elements, such as notes, clefs, time signatures, and other notations, before reasoning about their relationships to answer specific questions. This dual process presents two major challenges: (1) accurately identifying musical elements in the image, and (2) reasoning about these elements to generate appropriate answers.
Effectiveness of Reasoning.
Table 2: Evaluation results of frontier models for textual and visual QA on the SSMR-bench. The table presents the performance of various models across different sheet music reasoning tasks, including Rhythm, Chord, Interval, and Scale, with the overall performance summarized for both textual and visual QA.
Models Synthesize Sheet Music Reasoning Benchmark
@@
Qwen3-8B $54.00$ $60.50$ $50.75$ $66.25$ $57.88$
Qwen2.5-32B-Instruct $54.50$ $47.00$ $53.75$ $49.75$ $51.25$
@@
Llama3.1-70B-Instruct $39.75$ $41.00$ $47.00$ $48.00$ $43.94$
Llama3.1-8B-Instruct $23.50$ $19.00$ $18.00$ $23.25$ $20.94$

### v2

# Towards an AI Musician: Synthesizing Sheet Music Problems for Musical Reasoning

@@ Abstract
Enhancing the ability of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) to interpret sheet music is a crucial step toward building AI musicians. However, current research lacks both evaluation benchmarks and training data for sheet music reasoning. Inspired by mathematics, where simple operations yield infinite verifiable problems, we introduce a novel approach that treats core music theory rules, such as those governing beats and intervals, as programmatic functions to systematically synthesize a vast and diverse corpus of sheet music reasoning problems. This approach allows us to introduce a data synthesis framework that generates verifiable sheet music questions in both textual and visual modalities, leading to the Synthetic Sheet Music Reasoning Benchmark (SSMR-Bench) and a complementary training set. Evaluation results on SSMR-Bench highlight the key role reasoning plays in interpreting sheet music, while also pointing out the ongoing challenges in understanding sheet music in a visual format. By leveraging synthetic data for RLVR, all models show significant improvements on the SSMR-Bench. Additionally, they also demonstrate considerable advancements on previously established human-crafted benchmarks, such as MusicTheoryBench and the music subset of MMMU. Finally, our results show that the enhanced reasoning ability can also facilitate music composition.
@@ 1 Introduction
Figure 1: Comparison of Sheet Music Reasoning QA, Knowledge QA, and Music Recognition.
@@
Recent advancements in Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) have inspired researchers to explore the potential of developing AI musicians (Qu et al., 2025; Bradshaw & Colton, 2025; Wang et al., 2024). Given that sheet music is the universal language of musicians, the ability to read and interpret it is an essential step for AI musicians (Yuan et al., 2024; Wang et al., 2025). We term this capability sheet music reasoning. As illustrated in Figure 1, sheet music reasoning differs fundamentally from Music Knowledge QA (Li et al., 2024), which evaluates memorized knowledge, and from sheet music recognition (Chen et al., 2025a), which focuses on identifying notation from images. Instead, sheet music reasoning requires applying learned musical knowledge to actively interpret and analyze sheet music. This task requires not only accurate recognition of musical symbols but also a nuanced grasp of their interactions within broader musical structures.
However, research on sheet music reasoning is still limited, with few standardized benchmarks. For instance, ChatMusician (Yuan et al., 2024) introduces MusicTheoryBench, 367 human-designed questions partly targeting sheet music reasoning in ABC notation. MMMU (Yue et al., 2024) evaluates MLLMs on large-scale, college-level tasks, including some sheet music reasoning in image format. Yet both datasets are fully human-crafted, limiting scalability for training.
@@
Inspired by mathematics, where simple operations yield limitless verifiable problems, we propose to leverage the rules of music theory to programmatically generate an endless supply of verifiable reasoning problems about sheet music. Such a resource enables robust evaluation and supports training with verifiable rewards (DeepSeek-AI et al., 2025) to improve musical reasoning. We introduce a novel data synthesis framework that produces questions in both textual and visual staff notation, organized into nine templates across Rhythm, Chord, Interval, and Scale. Using this framework, we construct the Synthetic Sheet Music Reasoning Benchmark (SSMR-Bench), comprising 1,600 textual and 1,600 visual QA pairs for evaluation, plus 8,000 pairs per modality for training.
@@
We evaluate SSMR-Bench using a range of state-of-the-art LLMs and MLLMs. For the textual QA task, which represents sheet music in ABC notation, Qwen3-32B demonstrates performance on par with Deepseek-V3, while Deepseek-R1 exceeds Deepseek-V3 by 18.69% overall. This highlights the model’s enhanced reasoning capability in understanding sheet music. For the visual QA task, Gemini2.5-Pro achieves the best performance with an average score of 56.62%, surpassing Qwen2.5-VL-72B-Instruct, InternVL3-78B, and other models. However, this performance is considerably lower than Gemini2.5-Pro’s results on the same questions presented in textual format, highlighting the challenges of interpreting sheet music from images.
We train a range of models on our synthetic dataset using Group Relative Policy Optimization (GRPO,  Shao et al., 2024). Across different model families (Qwen, Yang et al., 2025 and Llama, Grattafiori et al., 2024), sizes (3B–8B), and modalities (text and image), all models achieve substantial improvements on SSMR-Bench. Notably, Qwen3-8B-Base improves by 13.06%, reaching performance comparable to GPT-4.1-mini, while Qwen2.5-VL-7B-Instruct exhibits stronger multimodal reasoning, surpassing Gemini2.5-Pro. Beyond the in-distribution benchmark, evaluation on external benchmarks such as MusicTheoryBench (Yuan et al., 2024) and MMMU Music (Yue et al., 2024), also demonstrates consistent improvement across all models. Interestingly, training on synthetic musical data also improves mathematical reasoning, underscoring the generality of the elicited capabilities. Moreover, our evaluations show that the enhanced reasoning ability leads to more rhythmically precise sheet music composition, with the trained Qwen3-8B-Base outperforming Qwen3-8B-Thinking and demonstrating greater musical coherence and accuracy.
@@
  • We are the first to leverage music theory rules to programmatically synthesize verifiable sheet music problems, using them as both an evaluation benchmark and a training set for Reinforcement Learning with Verifiable Rewards (RLVR).
@@ 2 Synthesize Music Reasoning Questions
2.1 Fundamentals of Sheet Music
Music is traditionally written on a staff, a set of five lines where visual symbols represent the core elements of music theory. The vertical position of a notehead indicates its pitch, while its shape (e.g., hollow or filled, with a stem or flag) defines its rhythm, or duration. These visual cues are organized into measures, separated by barlines, according to a time signature that dictates the underlying rhythmic pulse. While staff notation is intuitive for human musicians, its graphical nature is cumbersome for programmatic analysis. To bridge this gap, we utilize ABC notation, a text-based format that encodes the same musical information in a machine-readable way (Qu et al., 2025). In ABC, pitches are represented by letters (C, D, E), rhythmic values are specified with numerical modifiers, and measures are delineated by vertical bars (|). This format allows us to translate the abstract rules of music theory into parsable strings. Examples of ABC notation are provided in Appendix 6.
2.2 Data Synthesis Framework
Figure 2: Overview of synthesizing verifiable sheet music questions. The upper plot illustrates the pipeline. Experts define rule-based Question Template Classes that generate questions and answers from existing sheet music in ABC notation, resulting in textual QA problems. The sheet music is then converted into an image, allowing the creation of Visual Question Answering (VQA)-style questions. The process is fully programmatic, without LLMs. The lower plot presents examples of generated textual and visual QA tasks.
[…]
Figure 3: Algorithm for bar placement question synthesis, including bar removal, distractor generation, option shuffling, and rendering of textual and visual questions.
Our framework is designed to generate questions that test a model’s understanding of the fundamental music principles, from rhythmic calculation and chord identification to the analysis of intervals and scales. We outline the process of synthesizing sheet music questions based on music theory rules in Figure 2. Sheet music is first obtained in ABC notation. Then, the experts design specialized Question Template Classes derived from music theory rules, each capable of generating both questions and corresponding answers from the ABC representation. Applying these templates produces textual QA pairs. Subsequently, the sheet music in each question can be converted into images using abcm2ps and ImageMagick, enabling the construction of VQA questions. The entire pipeline is fully programmatic, without reliance on LLMs, ensuring controllability and interpretability.
The lower plot in Figure 2 presents a visual illustration of an example generated by the Bar Placement Class. In this case, the bar lines of the original sheet music are removed, and the resulting incomplete score serves as input. According to music theory, the missing bar lines can be reconstructed by reasoning from the time signature, requiring models to infer rhythmic grouping and meter to restore the bar-line structure. Once the textual QA question is obtained, the sheet music in textual modality can be converted into images to generate the corresponding VQA question. Figure 3 provides the pseudocode for the Bar Placement Class, demonstrating how it functions.
@@ 2.3 Data Construction
Table 1: Comparison of music-related QA datasets across different aspects, where checkmarks (✓) indicate presence and crosses (✗) indicate absence.
@@
MusicTheoryBench (Yuan et al., 2024)
SSMR-Bench (Ours)
The scalable synthesis framework we designed currently provides 9 types of questions. All questions are categorized into four categories: Rhythm, Chord, Interval, and Scale, which are the fundamental elements of sheet music. Detailed descriptions of the question types and categories can be found in Appendix 6. To solve these questions, LLMs and MLLMs need basic music knowledge, sheet music reading skills, and reasoning ability.
Through our data synthesis framework, we introduce the Synthetic Sheet Music Reasoning Benchmark (SSMR-Bench), comprising 1,600 textual and 1,600 visual questions. The sheet music used to generate questions is sampled from MelodyHub (Wu et al., 2024), a large-scale dataset in ABC notation. The distribution of question types and categories within the benchmark is illustrated in Figure 4. Furthermore, we offer a comprehensive training set, consisting of 8,000 textual and 8,000 visual questions. A comparative analysis between our framework and prior related benchmarks is presented in Table 1. Our framework produces both visual and textual modality questions for sheet music problems and constructs a verifiable training set without human curation, thereby enabling scalable data generation.
2.4 Verifiable Music Training
Following the RLVR paradigm, we employ the Group Relative Policy Optimization (GPRO, Shao et al., 2024) algorithm for RL training. GPRO approximates the advantage of a trajectory by normalizing its reward against the mean and standard deviation of rewards from $N$ sampled solutions for a given query:
$A_{i}=\frac{R(\tau_{i})-\text{mean}(\{R(\tau_{j})\mid j=1,\ldots,N\})}{\text{std}(\{R(\tau_{j})\mid j=1,\ldots,N\})},$ (1)
where $R(\tau_{i})$ is the reward for trajectory $\tau_{i}$. The rewards are calculated over $N$ trajectories sampled from the previous policy $\pi_{\theta_{\text{old}}}$. This relative scoring helps the model learn which outputs perform better or worse than the group’s average. Then, the GPRO objective is given by:
$\begin{gathered}\mathcal{J}_{\text{GRPO}}(\theta)=\frac{1}{\sum_{i=1}^{N}|\tau_{i}|}\sum_{i=1}^{N}\sum_{t=1}^{|\tau_{i}|}\text{CLIP}(r_{i,t}(\theta),A_{i},\epsilon)-\beta\cdot\mathbb{D}_{\text{KL}}[\pi_{\theta}\|\pi_{\text{ref}}].\end{gathered}$ (2)
Here, $r_{i,t}(\theta)={\pi_{\theta}(\tau_{i,t}|q,\tau_{i,<t})}/{\pi_{\theta_{\text{old}}}(\tau_{i,t}|q,\tau_{i,<t})}$} is the importance sampling ratio, adjusting for the fact that the trajectories were generated by the old policy $\pi_{\theta_{\text{old}}}$, ensuring unbiased policy updates.
@@ 3.1 Experimental Setup
Implementation Details.
We employ multiple LLMs and MLLMs for reinforcement learning on our synthetic data. For the textual modality, we choose Qwen3-4B-Base, Qwen3-8B-Base, and Llama3.1-8B-IT (with “IT” referring to “Instruct”). For the visual modality, we select Qwen2.5-VL-7B-Instruct. Additionally, we include a variant of Qwen3-8B-Base trained on 8,000 math problems sampled from OpenR1 (Hugging Face, 2025) to compare the effectiveness of synthetic music reasoning data. Training is conducted using the GRPO algorithm with a batch size of 128, a rollout number of 8, and an update batch size of 64 over 300 training steps. The KL loss and entropy loss are removed. The reward function is based solely on accuracy, without incorporating format-related signals. The training template is provided in Appendix B.
@@ 3.2 Results on Synthetic Music Reasoning Benchmark
Table 2: Evaluation results of frontier models for textual and visual QA on the SSMR-bench. The table presents the performance of various models across different sheet music reasoning tasks, including Rhythm, Chord, Interval, and Scale, with the overall performance summarized for both textual and visual QA. “IT” is an abbreviation for “Instruct”.
Models Thinking Synthetic Sheet Music Reasoning Benchmark
@@
Qwen3-8B-Thinking $54.00$ $60.50$ $50.75$ $66.25$ $57.88$
Qwen2.5-32B-IT $54.50$ $47.00$ $53.75$ $49.75$ $51.25$
@@
Llama3.1-70B-IT $39.75$ $41.00$ $47.00$ $48.00$ $43.94$
Llama3.1-8B-IT $23.50$ $19.00$ $18.00$ $23.25$ $20.94$
@@
Textual Modality Performance Comparison.
The results for textual QA clearly show that LLMs, such as GPT-5, outperform other models across all categories. Specifically, GPT-5 attains the highest overall accuracy at 97.00%, followed by Gemini-2.5-Pro at 96.10%, DeepSeek-R1 at 93.63%, and GPT-4.1 at 79.81%. In contrast, models like Qwen2.5-32B-IT and Llama3.1-70B-IT exhibit lower overall performance, with scores of 51.25% and 43.94%, respectively. Llama3.1-8B-IT shows the lowest performance, with an overall score of 20.94%.
Performance of MLLMs on Visual Modality.
Cross-Modality Performance Analysis.
The lower performance of MLLMs in the visual modality can be attributed to the unique challenges posed by music element recognition. For Gemini-2.5-Pro, accuracy reaches 96.19% when sheet music is presented in textual modality, compared to 55.44% when presented in visual modality. Unlike textual data, which is relatively machine-friendly, musical staff images contain intricate visual patterns. To interpret these images, MLLMs must first recognize diverse musical elements, such as notes, clefs, time signatures, and other notations, before reasoning about their relationships to answer specific questions. This dual process presents two major challenges: (1) accurately identifying musical elements in the image, and (2) reasoning about these elements to generate appropriate answers.
Effect of Reasoning on Performance.